OpenStudy (anonymous):

x^2-y^2/x-y * 2xy/x^2+xy someone help pleaseee.!

OpenStudy (akashdeepdeb):

\[\frac{(x^2-y^2)(2xy)}{(x-y)(x^2+xy)}\] There is an algebraic identity: \(x^2-y^2 = (x+y)(x-y)\) From \(x^2+xy\) take '\(x\)' common, you'll have: \(x(x+y)\) Writing it like that, we have: \[\frac{(x+y)(x-y)(2xy)}{(x-y)(x+y)x}\] Can you proceed now? :)

OpenStudy (anonymous):

i got 2x2y • (x + y)2

OpenStudy (akashdeepdeb):

How exactly? If you see, \(x(x+y)(x-y)\) are all present in the numerator and in the denominator. They cancel each other out don't they? Look at it carefully, once more! :)